OpenStudy (anonymous):

One-half the sum of four consecutive integers is 37. Find the integers.

OpenStudy (anonymous):

(17,18,19,20) because (17+18+19+20)/2=37.

OpenStudy (anonymous):

To get this answer, you figure out that the sum of the four integers you need is 2*37=74. 74/4=18.5. Therefore, the integers are 17, 18, 19, 20

OpenStudy (anonymous):

but 20 is even

OpenStudy (anonymous):

(x+(x+1)+(x+2)+(x+3))/2 =37

OpenStudy (anonymous):

Can you solve the rest?

OpenStudy (anonymous):

I said one of the intergers was x so since there are increasing each one must increase by 1 more than the last value.

OpenStudy (anonymous):

i hear you but i dont know how to find the value of x?

OpenStudy (anonymous):

Well add all the x's up.

OpenStudy (anonymous):

i got 4x and 5

OpenStudy (anonymous):